# Logarithmic Sobolev inequality in manifolds with nonnegative curvature via the ABP method

### v2

@@ 1. Introduction
$\int_{\mathbb{R}^{n}}f^{2}\log f^{2}\,d\gamma-\Big(\int_{\mathbb{R}^{n}}f^{2}\,d\gamma\Big)\log\Big(\int_{\mathbb{R}^{n}}f^{2}\,d\gamma\Big)\leq 2\int_{\mathbb{R}^{n}}|Df|^{2}\,d\gamma,$
@@
$\int_{\mathbb{R}^{n}}f^{2}\log f^{2}\,dx\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\int_{\mathbb{R}^{n}}|Df|^{2}\,dx\Big),$
@@
$\int_{\mathbb{R}^{n}}|f|^{p}\log|f|^{p}\,dx\leq\frac{n}{p}\log\Big(\mathrm{L}(n,p)\int_{\mathbb{R}^{n}}|Df|^{p}\,dx\Big),$
@@
$\mathrm{L}(n,p)=\left\{\begin{array}[]{ll}n^{-1}\omega_{n}^{-\frac{1}{n}}&\text{if }p=1,\\ \frac{p}{n}\Big(\frac{p-1}{e}\Big)^{p-1}\Big(\omega_{n}\Gamma\big(\frac{n}{q}+1\big)\Big)^{-\frac{p}{n}}&\text{if }p>1.\end{array}\right$.
@@
Recently, by generalizing the optimal transport approach developed by Cordero-Erausquin, Nazaret and Villani [10], Kristály [17] obtained the optimal $L^{p}$ logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature. In this paper, inspired by the ABP method developed by Brendle [5], our first goal is to provide a ABP-based proof of this inequality.
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with smooth boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$. Then
$\int_{\Omega}f\log f\,d\mathrm{vol}\leq n\log\bigg(n^{-1}\omega_{n}^{-\frac{1}{n}}\theta^{-\frac{1}{n}}\Big(\int_{\Omega}|Df|\,d\mathrm{vol}+\int_{\partial\Omega}f\,da\Big)\bigg),$
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with smooth boundary $\partial\Omega$, and let $f$ be a positive smooth function in $\Omega\setminus\partial\Omega$ that vanishes on $\partial\Omega$. For $p>1$, we have
@@
Let $(M^{n+m},g)$, with $n\geq 2$ and $m\geq 1$, be a complete noncompact Riemannian manifold with nonnegative sectional curvature and Euclidean volume growth. Let $\Sigma^{n}$ be a compact submanifold of $M$, possibly with boundary, and let $f$ be a positive smooth function in $\Sigma\setminus\partial\Sigma$ that vanishes on $\partial\Sigma$. Then
$\int_{\Sigma}f^{2}\log f^{2}\,d\mathrm{vol}\leq\frac{n}{2}\log\Big(\frac{2}{\pi en}\theta^{-\frac{2}{n}}\int_{\Sigma}|D^{\Sigma}f|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\Big).$
@@ 2. The case of domains
Throughout this section, we assume that $(M^{n},g)$, for $n\geq 2$, is a complete noncompact manifold with nonnegative Ricci curvature, and we fix a base point $o\in M$. Let $\Omega\subset M$ be a compact domain with smooth boundary $\partial\Omega$. Let $p\geq 1$ and let $f$ be a smooth positive function in $\Omega\setminus\partial\Omega$. In the case $p>1$, we assume that $f$ vanishes on $\partial\Omega$.
without loss of generality, we assume hereafter that $\Omega$ is connected. Let $\alpha>0$ be a given parameter and $r_{o}=d(o,\Omega)$. By scaling, we may assume that
(1) $\left\{\begin{aligned} &\int_{\Omega}f^{p}\log f^{p}\,d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Df|^{p}\,d\text{vol}=\frac{1}{\alpha}\int_{\partial\Omega}f\,da,&p=1,\\ &\int_{\Omega}f^{p}\log f^{p}d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Df|^{p}d\text{vol}=0,&p>1.\end{aligned}\right$.
Since $\Omega$ is connected, we can find a solution of the following Neumann boundary problem
(2) $\left\{\begin{aligned} &\mathrm{div}(f^{p}Du)=f^{p}\log f^{p}-\frac{1}{\alpha}|Df|^{p},&\text{ in }\Omega\setminus\partial\Omega,\\ &\langle Du,\nu\rangle=\frac{r^{p-1}}{\alpha},&\text{ on }\partial\Omega,\end{aligned}\right$.
where $\nu$ is the outward unit normal vector field and $r>1+\alpha r_{o}$ is an arbitrary constant. By standard elliptic regularity theory, we know that $u\in C^{2,\gamma}$ for each $0<\gamma<1$.
As in [5], we set $U:=\{x\in\Omega\setminus\partial\Omega:|Du(x)|<\frac{r^{p-1}}{\alpha}\}$ and
@@
$\Omega_{r}:=\{\bar{x}\in M:d(x,\bar{x})<\frac{r^{p}}{\alpha},\quad\forall x\in\Omega\}$
@@
$|\det D\Phi_{r}(x)|\leq(1+r\frac{\Delta u(x)}{n})^{n}$
@@
$\left\{\begin{aligned} &|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}f(x),&p=1,\\ &e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{|d(x,\Phi_{r}(x))|^{q}}{r^{q}}}|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}f(x)^{p},&p>1.\end{aligned}\right$.
@@
$\mathrm{div}(f^{p}Du)$ $=f^{p}\Delta u+p\langle f^{p-1}Df,Du\rangle$
$=f^{p}\log f^{p}-\frac{1}{\alpha}|Df|^{p}.$
@@
$\Delta u$ $=\log f-\frac{1}{\alpha}\frac{|Df|}{f}-\langle\frac{Df}{f},Du\rangle$
$\leq\log f.$
@@
$|\det D\Phi_{r}(x)|$ $\leq\Big(\frac{1}{r}+\frac{\Delta u(x)}{n}\Big)^{n}r^{n}$
$\leq r^{n}e^{\frac{n}{r}+\Delta u-n}$
$=r^{n}e^{\frac{n}{r}-n}f(x).$
@@
$\Delta u$ $=\log f^{p}-\frac{|Df|^{p}}{\alpha f^{p}}-p\langle\frac{Df}{f},Du\rangle$
$\leq\log f^{p}+\frac{p}{q}\alpha^{\frac{q}{p}}|Du|^{q}.$
@@
$|\det D\Phi_{r}(x)|$ $\leq\Big(\frac{1}{r}+\frac{\Delta u(x)}{n}\Big)^{n}r^{n}$
@@
$\leq r^{n}e^{\frac{n}{r}-n}e^{\frac{p}{q}\alpha^{\frac{q}{p}}|Du(x)|^{q}}f(x)^{p},$
@@
$e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{|d(x,\Phi_{r}(x))|^{q}}{r^{q}}}|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}f(x)^{p}.$
@@
Using the triangle inequality, Lemma 2.1 and Lemma 2.3, the formula for change of variables in multiple integrals implies
@@
$\leq|\{\bar{x}\in M:d(x,\bar{x})<\frac{r}{\alpha},\,\forall x\in\Omega\}|$
@@
$\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}f(x)\,d\text{vol}(x)$
for all $r>1+\alpha r_{o}$. Dividing the above formula by $r^{n}$ and let $r\to\infty$, we obtain
$\theta\omega_{n}\alpha^{-n}\leq e^{-n}\int_{\Omega}f\,d\text{vol},$
@@
$n+\log(\theta\omega_{n}\alpha^{-n})\leq\log\bigg(\int_{\Omega}f\,d\text{vol}\bigg).$
@@
(3) $\int_{\Omega}f(\log f+n+\log(\theta\omega_{n}\alpha^{-n}))\,d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Df|\,d\text{vol}-\frac{1}{\alpha}\int_{\partial\Omega}f\,da$
$\leq\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big)\log\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big)$
@@
$\int_{\Omega}f\log f\,d\text{vol}\leq-n-\log(\theta\omega_{n}\alpha^{-n})+\frac{1}{\alpha}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big).$
By minimizing the right hand side with respect to $\alpha>0$, one can get
$\alpha=\frac{1}{n}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big).$
@@
$\int_{\Omega}f\log f\,d\text{vol}\leq n\log\bigg(n^{-1}\omega_{n}^{-\frac{1}{n}}\theta^{-\frac{1}{n}}\Big(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\Big)\bigg).$
@@
Let $p>1$. By the triangle inequality, Lemma 2.1 and Lemma 2.3, we have
@@
$\leq\int_{\Omega}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(d(x,\Phi_{r}(x)))^{q}}{r^{q}}}|\det D\Phi_{r}(x)|1_{A_{r}}(x)\,d\mathrm{vol}(x)$
$\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}f(x)^{p}\,d\mathrm{vol}(x).$
@@
$\int_{0}^{\frac{r^{p}}{\alpha}-r_{o}}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(r_{o}+\rho)^{q}}{r^{q}}}n\omega_{n}\theta\rho^{n-1}\,d\rho\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}f(x)^{p}\,d\mathrm{vol}(x).$
Dividing the above formula by $r^{n}$ and let $r\to\infty$, it follows from (A.1) in [1] that
$\theta e^{n}\bigg(\Big(\frac{p}{q}\alpha^{\frac{q}{p}}\Big)^{-\frac{n}{q}}\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\bigg)$
$=\lim_{r\to\infty}r^{-n}e^{n-\frac{n}{r}}\int_{0}^{\frac{r^{p}}{\alpha}-r_{o}}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(r_{o}+\rho)^{q}}{r^{q}}}n\omega_{n}\theta\rho^{n-1}\,d\rho$
$\leq\int_{\Omega}f(x)^{p}\,d\mathrm{vol}(x).$
Combining this inequality with (1) gives
(4) $\int_{\Omega}f^{p}\bigg(\log f^{p}+n+\log\theta-\frac{n}{q}\log\Big(\frac{p}{q}\alpha^{\frac{q}{p}}\Big)+\log\Big(\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\Big)\bigg)\,d\mathrm{vol}$
$\leq\frac{1}{\alpha}\int_{\Omega}|Df|^{p}\,d\mathrm{vol}+\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big)\log\Big(\int_{\Omega}f^{p}\,d\mathrm{vol}\Big).$

### v3

@@ 1. Introduction
$\int_{\mathbb{R}^{n}}f^{2}\log f^{2}\,d\gamma-\bigg(\int_{\mathbb{R}^{n}}f^{2}\,d\gamma\bigg)\log\bigg(\int_{\mathbb{R}^{n}}f^{2}\,d\gamma\bigg)\leq 2\int_{\mathbb{R}^{n}}|Df|^{2}\,d\gamma,$
@@
$\int_{\mathbb{R}^{n}}f^{2}\log f^{2}\,dx\leq\frac{n}{2}\log\bigg(\frac{2}{\pi en}\int_{\mathbb{R}^{n}}|Df|^{2}\,dx\bigg),$
@@
$\int_{\mathbb{R}^{n}}|f|^{p}\log|f|^{p}\,dx\leq\frac{n}{p}\log\bigg(\mathrm{L}(n,p)\int_{\mathbb{R}^{n}}|Df|^{p}\,dx\bigg),$
@@
$\mathrm{L}(n,p)=\left\{\begin{array}[]{ll}n^{-1}\omega_{n}^{-\frac{1}{n}}&\text{if }p=1,\\ \frac{p}{n}\Big(\frac{p-1}{e}\Big)^{p-1}\Big(\omega_{n}\Gamma\Big(\frac{n}{q}+1\Big)\Big)^{-\frac{p}{n}}&\text{if }p>1,\end{array}\right$.
@@
Recently, by generalizing the optimal transport approach developed by Cordero-Erausquin, Nazaret and Villani [10], Kristály [17] obtained the optimal $L^{p}(p\geq 1)$ logarithmic Sobolev inequality for manifolds with nonnegative Ricci curvature. In this paper, inspired by the ABP method developed by Brendle [5], our first goal is to provide a ABP-based proof of this inequality.
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with smooth boundary $\partial\Omega$, and let $f$ be a nonnegative smooth function on $\Omega$ with $\int_{\Omega}f\,d\mathrm{vol}=1$. Then
$\int_{\Omega}f\log f\,d\mathrm{vol}\leq n\log\bigg(n^{-1}\omega_{n}^{-\frac{1}{n}}\theta^{-\frac{1}{n}}\bigg(\int_{\Omega}|Df|\,d\mathrm{vol}+\int_{\partial\Omega}f\,da\bigg)\bigg),$
@@
Let $(M^{n},g)$, with $n\geq 2$, be a complete noncompact Riemannian manifold with nonnegative Ricci curvature and Euclidean volume growth. Let $\Omega$ be a compact domain in $M$ with smooth boundary $\partial\Omega$, and let $f$ be a nonnegative smooth function on $\Omega$ vanishing on $\partial\Omega$ such that $\int_{\Omega}f^{p}\,d\mathrm{vol}=1$ . For $p>1$, we have
@@
Let $(M^{n+m},g)$, with $n\geq 2$ and $m\geq 1$, be a complete noncompact Riemannian manifold with nonnegative sectional curvature and Euclidean volume growth. Let $\Sigma^{n}$ be a compact submanifold of $M$, possibly with boundary, and let $f$ be a nonnegative smooth function on $\Sigma$ vanishing on $\partial\Sigma$ with $\int_{\Sigma}f^{2}\,d\mathrm{vol}=1$. Then
$\int_{\Sigma}f^{2}\log f^{2}\,d\mathrm{vol}\leq\frac{n}{2}\log\bigg(\frac{2}{\pi en}\theta^{-\frac{2}{n}}\int_{\Sigma}|D^{\Sigma}f|^{2}+\frac{1}{4}f^{2}|H|^{2}\,d\mathrm{vol}\bigg),$
@@ 2. The case of domains
Throughout this section, we assume that $(M^{n},g)$, for $n\geq 2$, is a complete noncompact manifold with nonnegative Ricci curvature, and we fix a base point $o\in M$. Let $\Omega\subset M$ be a compact domain with smooth boundary $\partial\Omega$. Let $p\geq 1$ and let $f$ be a smooth nonnegative function on $\Omega$. In the case $p>1$, we assume that $f$ vanishes on $\partial\Omega$.
without loss of generality, we assume hereafter that $\Omega$ is connected. Let $\alpha>0$ be a given parameter, $r_{o}=\max\{d(o,x):x\in\Omega\}$ and $r>1+\alpha r_{0}$ is an arbitrary constant. By scaling, we may assume that
(1) $\int_{\Omega}h^{p}\log h^{p}\,d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Dh|^{p}\,d\text{vol}=\frac{r^{p-1}}{\alpha}\int_{\partial\Omega}h^{p}\,da,$
where $h=c(f+1/r)$ for some $c>0$. Since $\Omega$ is connected, we can find a solution of the following Neumann boundary problem
(2) $\left\{\begin{aligned} &\mathrm{div}(h^{p}Du)=h^{p}\log h^{p}-\frac{1}{\alpha}|Dh|^{p},&\text{ on }\Omega,\\ &\langle Du,\nu\rangle=\frac{r^{p-1}}{\alpha},&\text{ on }\partial\Omega,\end{aligned}\right$.
where $\nu$ is the outward unit normal vector field. By standard elliptic regularity theory, we know that $u\in C^{2,\gamma}$ for each $0<\gamma<1$.
As in [5], we set $U:=\{x\in\Omega\setminus\partial\Omega:|Du(x)|<r^{p-1}/\alpha\}$ and
@@
$\Omega_{r}:=\{\bar{x}\in M:d(x,\bar{x})<r^{p}/\alpha,\,\forall x\in\Omega\}$
@@
$|\det D\Phi_{r}(x)|\leq\bigg(1+r\frac{\Delta u(x)}{n}\bigg)^{n}$
@@
$\left\{\begin{aligned} &|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}h(x),&p=1,\\ &e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{d(x,\Phi_{r}(x))^{q}}{r^{q}}}|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}h(x)^{p},&p>1,\end{aligned}\right$.
@@
$\mathrm{div}(h^{p}Du)$ $=h^{p}\Delta u+p\langle h^{p-1}Dh,Du\rangle$
$=h^{p}\log h^{p}-\frac{1}{\alpha}|Dh|^{p}.$
@@
$\Delta u$ $=\log h-\frac{|Dh|}{\alpha h}-\frac{\langle Dh,Du\rangle}{h}$
$\leq\log h.$
@@
$|\det D\Phi_{r}(x)|$ $\leq\bigg(\frac{1}{r}+\frac{\Delta u(x)}{n}\bigg)^{n}r^{n}$
$\leq r^{n}e^{\frac{n}{r}+\Delta u(x)-n}$
$=r^{n}e^{\frac{n}{r}-n}h(x).$
@@
$\Delta u$ $=\log h^{p}-\frac{|Dh|^{p}}{\alpha h^{p}}-p\frac{\langle Dh,Du\rangle}{h}$
$\leq\log h^{p}+\frac{p}{q}\alpha^{\frac{q}{p}}|Du|^{q}.$
@@
$|\det D\Phi_{r}(x)|$ $\leq\bigg(\frac{1}{r}+\frac{\Delta u(x)}{n}\bigg)^{n}r^{n}$
@@
$\leq r^{n}e^{\frac{n}{r}-n}e^{\frac{p}{q}\alpha^{\frac{q}{p}}|Du(x)|^{q}}h(x)^{p},$
@@
$e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{d(x,\Phi_{r}(x))^{q}}{r^{q}}}|\det D\Phi_{r}(x)|\leq r^{n}e^{\frac{n}{r}-n}h(x)^{p}.$
@@
Let $p=1$. Using the triangle inequality, Lemma 2.1 and Lemma 2.3, the formula for change of variables in multiple integrals implies
@@
$\leq|\{\bar{x}\in M:d(x,\bar{x})<r/\alpha,\,\forall x\in\Omega\}|$
@@
$\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}h(x)\,d\text{vol}(x)$
for all $r>1+\alpha r_{o}$. Dividing the above formula by $r^{n}$, we obtain
$\theta\omega_{n}(\alpha^{-1}-r_{o}/r)^{n}\leq e^{\frac{n}{r}-n}\int_{\Omega}h\,d\text{vol},$
@@
$n-n/r+\log(\theta\omega_{n}(\alpha^{-1}-r_{o}/r)^{n})\leq\log\bigg(\int_{\Omega}h\,d\text{vol}\bigg).$
@@
$\int_{\Omega}h(\log h+n-n/r+\log(\theta\omega_{n}(\alpha^{-1}-r_{o}/r)^{n}))\,d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Dh|\,d\text{vol}$
$\leq\bigg(\int_{\Omega}h\,d\mathrm{vol}\bigg)\log\bigg(\int_{\Omega}h\,d\mathrm{vol}\bigg)+\frac{1}{\alpha}\int_{\partial\Omega}h\,da.$
Taking $h=f+\frac{1}{r}$ with $r\to\infty$, one can get
(3) $\int_{\Omega}f(\log f+n+\log(\theta\omega_{n}\alpha^{-n}))\,d\text{vol}-\frac{1}{\alpha}\int_{\Omega}|Df|\,d\text{vol}$
$\leq\bigg(\int_{\Omega}f\,d\mathrm{vol}\bigg)\log\bigg(\int_{\Omega}f\,d\mathrm{vol}\bigg)+\frac{1}{\alpha}\int_{\partial\Omega}f\,da.$
@@
$\int_{\Omega}f\log f\,d\text{vol}\leq-n-\log(\theta\omega_{n}\alpha^{-n})+\frac{1}{\alpha}\bigg(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\bigg).$
By minimizing the right hand side with respect to $\alpha>0$, one has
$\alpha=\frac{1}{n}\bigg(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\bigg).$
@@
$\int_{\Omega}f\log f\,d\text{vol}\leq n\log\bigg(n^{-1}\omega_{n}^{-\frac{1}{n}}\theta^{-\frac{1}{n}}\bigg(\int_{\Omega}|Df|\,d\text{vol}+\int_{\partial\Omega}f\,da\bigg)\bigg).$
@@
Let $p>1$. By the triangle inequality, Lemma 2.1 and Lemma 2.3, we find
@@
$\leq\int_{\Omega}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{d(x,\Phi_{r}(x))^{q}}{r^{q}}}|\det D\Phi_{r}(x)|1_{A_{r}}(x)\,d\mathrm{vol}(x)$
$\leq r^{n}e^{\frac{n}{r}-n}\int_{\Omega}h(x)^{p}\,d\mathrm{vol}(x).$
@@
$r^{-n}e^{n-\frac{n}{r}}\int_{0}^{\frac{r^{p}}{\alpha}-r_{o}}e^{-\frac{p}{q}\alpha^{\frac{q}{p}}\frac{(r_{o}+\rho)^{q}}{r^{q}}}n\omega_{n}\theta\rho^{n-1}\,d\rho\leq\int_{\Omega}h(x)^{p}\,d\mathrm{vol}(x).$
Combining this inequality with (1) and letting $h=f+\frac{1}{r}$ as $r\to\infty$, it follows from (A.1) in [1] that
(4) $\int_{\Omega}f^{p}\bigg(\log f^{p}+n+\log\theta-\frac{n}{q}\log\bigg(\frac{p}{q}\alpha^{\frac{q}{p}}\bigg)+\log\bigg(\omega_{n}\Gamma\bigg(\frac{n}{q}+1\bigg)\bigg)\bigg)\,d\mathrm{vol}$
$\leq\frac{1}{\alpha}\int_{\Omega}|Df|^{p}\,d\mathrm{vol}+\bigg(\int_{\Omega}f^{p}\,d\mathrm{vol}\bigg)\log\bigg(\int_{\Omega}f^{p}\,d\mathrm{vol}\bigg).$